OpenStudy (lilai3):

80% of all CA drivers wear seat belts. If 3 drivers are pulled over, what's the probability that all would be wearing their seat belts? Write as a percent to the nearest tenth and explain. Thx!

OpenStudy (kinggeorge):

Assuming the 3 drivers are random, there is a 8/10 chance that each individual is wearing a seat belt. If you know that, do you know how to find the probability that 3 random drivers are wearing a seatbelt?

OpenStudy (lilai3):

uh... i don't get it....?

OpenStudy (kinggeorge):

If you have a .8 (=8/10) chance of something happening once, you have a \[\left({8\over10}\right)^3=(.8)^3 = .512 =51.2\%\]Of it happening 3 times.

OpenStudy (lilai3):

so the answer is...51.2%?

OpenStudy (kinggeorge):

Yes. Just multiply the original probability with itself 3 times since you are pulling over 3 random drivers.

OpenStudy (lilai3):

oh..dat's how u do it.... thank u....(dang it.. i;'m always having trouble with these.. so ya..... i appreciate dat ur giving time to answer my questions! ur da best so far! ;)

OpenStudy (lilai3):

and is the original probability 80%?

OpenStudy (kinggeorge):

You're very welcome. :) And yes, the original probability is 80%

OpenStudy (lilai3):

thx